OpenStudy (anonymous):

A student has test scores of 72, 80, 65, 78, and 60. If the final exam counts counts as one third of the final grade what scores must the student receive in order to have a final grade of 76?

OpenStudy (anonymous):

an average of 126.6 should be scored

OpenStudy (anonymous):

Final = 76 Average Test Score = Sum of tests/number of tests = (72+80+65+78+60)/5 Final = 2/3 Average test Score + 1/3 Final Exam Score = 76 Just plug n chug. I don't feel like doing it.